# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ 5.5.3 Proposed models
As for the proposed models, 18 out of 39 belong to the pure convolutional GNN paradigm, followed by 11 attentional GNNs, 9 hybrid convolutional-attention architectures, and 1 recurrent GNN. Regarding the definition of the graph structure, most of the papers discuss it in detail. Half of the papers are based on a pre-defined adjacency matrix, and among the remaining papers that propose models that learn the adjacency matrix on their own, a further half is based on a pre-initialization of the matrix, as for example in Ref. [art_108]. Here, more than in any other group, there are multiple ways to define the adjacency matrix, taking into account either the structure of the road network, the actual flow of people, or a combination of different information. More in detail, some models use the road connectivity ([art_104], [art_120], [art_123], [art_135]), others their spatial distance ([art_93], [art_96], [art_99], [art_110], [art_128], [art_132]), and some others incorporate extra information about the so-called points of interest (such as Refs. [art_94], [art_96]). Other models instead consider some similarity measures between the time series ([art_97], [art_117], [art_122], [art_123]), such as the correlation coefficient or cosine similarity. Still others directly use the number of people traveling on the road at a given time step ([art_94], [art_102], [art_129], [art_133]).
@@
The two most common loss functions used to train the models are the MSE and the MAE. Few papers consider a combination of loss functions, especially when dealing with a model that has to learn the graph structure by itself (Refs. [art_95], [art_126], [art_130]). Not all papers specify the language or library used for the code. Among those that do, Python is the most common language, with the majority using PyTorch, while only a few researchers use TensorFlow. In addition, a few papers provide a link to the source code for the proposed model. The papers that provide such links are listed in Tab. LABEL:t:link_codici_traffico.
Table 17: List of source codes of the ”Mobility” models in the review.
@@
Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [art_108] https://github.com/jinguangyi-uto-DSTSGN
@@ 5.5.4 Benchmark models
Tab. LABEL:t:benchmark_models_traffico lists the benchmark models used in the papers in the ”Mobility” group. The benchmarks are categorized into mathematical and statistical methods, classical machine learning models, and GNN models, and only those appearing in more than two papers are included.
Table 18: List of benchmark models in the ”Mobility” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_93], [art_94], [art_95], [art_96], [art_98], [art_100], [art_101], [art_103], [art_106], [art_107], [art_109], [art_116], [art_123], [art_126], [art_129], [art_132], [art_135]
Historical average (HA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_101], [art_102], [art_103], [art_109], [art_110], [art_120], [art_123], [art_125], [art_131], [art_132], [art_135]
Vector autoregression (VAR) [art_99], [art_100], [art_126], [art_131], [art_132]
Classical machine learning methods Feed-forward neural network (FNN) [art_93], [art_97], [art_100], [art_105], [art_114], [art_120], [art_121], [art_124], [art_126]
Gated recurrent unit (GRU) [art_98], [art_99], [art_101], [art_106], [art_121], [art_123], [art_124], [3], [art_131], [art_132], [art_135]
Long-short term memory network (LSTM) [art_93], [art_96], [art_98], [art_99], [art_100], [art_106], [art_107], [art_108], [art_109], [art_113], [art_115], [art_116], [art_120], [art_121], [art_122], [art_125], [art_126], [3], [art_130], [art_131], [art_132]
@@
Seq2Seq [Seq2seq] [art_105], [art_124]
Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [ST-Norm] [art_116], [art_117]
Support vector regression (SVR) [art_99], [art_100], [art_103], [art_107], [art_109], [art_119], [art_123], [art_129], [art_130], [art_135]
@@
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_95], [art_98], [art_100], [art_101], [art_110], [art_117], [art_124], [art_125], [art_130], [art_131]
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_93], [art_95], [art_98], [art_99], [art_100], [art_107], [art_108], [art_110], [art_113], [art_115], [art_117], [art_121], [art_125], [art_126], [3], [art_130], [art_131], [art_132]
@@
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_93], [art_94], [art_95], [art_96], [art_100], [art_101], [art_103], [art_107], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_124], [art_125], [art_126], [art_128], [art_129], [art_130], [art_131], [art_132]
Dynamic spatial-temporal aware graph neural network (DSTAGNN) [DSTAGNN] [art_95], [art_125]
@@
Graph convolutional network (GCN) [art_104], [art_135]
Graph multi-attention network (GMAN) [GMAN] [art_100], [art_109], [art_110], [art_116], [art_122], [art_128]
Graph WaveNet (GWN) [GWN] [art_93], [art_95], [art_100], [art_101], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_122], [art_124], [art_128]
Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [MRA-BGCN] [art_109], [art_116]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_100], [art_109], [art_110], [art_113], [art_116], [art_117], [art_122], [art_124], [art_128]
@@
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_98], [art_101], [art_108], [art_115], [art_117], [art_125], [art_130]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_93], [art_94], [art_95], [art_96], [art_99], [art_100], [art_101], [art_106], [art_107], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_121], [art_122], [art_124], [art_126], [3], [art_128], [art_129], [art_130], [art_131], [art_132]
@@
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_99], [art_108], [art_117], [art_125]
Deep-meta-learning based model (ST-MetaNet) [ST-MetaNet] [art_96], [art_100], [art_109], [art_116]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_98], [art_99], [art_100], [art_108], [art_115], [art_117], [art_125], [art_130], [art_131]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_103], [art_107], [art_113], [art_121], [art_123]
As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and historical average model (HA), which uses the average of past observations as the forecast for future values. Despite their apparent simplicity, these models are sometimes shown to be very powerful benchmarks, with an accuracy comparable to that of more complex models. As for the traditional machine learning models, LSTMs are widely recognized as a type of recurrent neural network capable of handling and long-term dependencies in traffic data. As for GNN-based models instead, the three most used models are, in order, the spatio-temporal graph convolutional network (STGCN) [STGCN], the diffusion convolutional recurrent neural network (DCRNN) [DCRNN], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], all developed specifically for the traffic forecasting problem.
The STGCN model was proposed by Yu et al. in 2018, and it integrates graph convolution and gated temporal convolution through spatio-temporal convolutional blocks. The source code is available at https://github.com/VeritasYin/STGCN_IJCAI-18. The DCRNN model was proposed by Li et al. in 2017, and it aims to capture the spatial dependency using bidirectional random walks on the graph, and the temporal dependency using the encoder-decoder architecture with scheduled sampling. The code can be accessed at https://github.com/liyaguang/DCRNN. Finally, the ASTGCN model was proposed by Guo et al. in 2019, and it models in parallel recent, daily-periodic and weekly-periodic dependencies as three independent components with a spatio-temporal attention mechanism to capture the dynamic spatio-temporal correlations, and a spatio-temporal convolution which simultaneously employs graph convolutions to capture the spatial patterns. The code is available at https://github.com/Davidham3/ASTGCN-2019-mxnet. For a more comprehensive list of source codes for the most dated benchmark GNN models in the traffic domain see Ref. [review_traffico_jiang].
@@ 5.5.5 Results
In all papers, the authors claim that the proposed GNN-based models are able to outperform the other benchmarks. However, there are a limited number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs. LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay.
For the and comparison and evaluation of the models, the most commonly used error metrics are MAE, RMSE, and MAPE. Few papers also use different metrics, such as the coefficient of determination $R^{2}$. However, despite the use of the same metrics and the availability of widely used datasets, comparing models is not trivial because different papers often focus on different time windows of a dataset.
@@
Table 19: Comparison of the average accuracy of the different models on the METR-LA dataset for 15, 30 and 60 minute time horizons, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
@@
The results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon. However, especially for longer time horizons, the Dynamic Graph Convolution LSTM Network (DyGCN-LSTM) [art_113] appears to be the most accurate one (excluding the ”Generic” models, among which the best model is the multichannel fusion graph neural network (MCFGNN) [art_86]).
@@
Table 20: Comparison of the average accuracy of the different models on the PEMS-BAY dataset for 15, 30 and 60 minute time horizons, expressed in terms of MAE, MAPE and RMSE. The smallest errors are underlined. Numbers from the papers.
@@
The results on the the PEMS-BAY dataset indicate that the dynamic graph convolutional recurrent network (DGCRN) [art_100] is the most accurate model for the 15-minute forecasting horizon, while the multi-view spatial-temporal graph neural network (MVST-GNN) [art_109] is the best model for the longer horizons.
@@
Table 21: Comparison of the average accuracy of the different models on the PeMS03 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
@@
Table 22: Comparison of the average accuracy of the different models on the PeMS04 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
@@
Table 23: Comparison of the average accuracy of the different models on the PeMS07 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The smallest errors are underlined. Numbers from the papers.
@@
DCRNN 25.30 11.66 38.58
@@
Table 24: Comparison of the average accuracy of the different models on the PeMS08 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the ”Generic” group. The smallest errors are underlined. Numbers from the papers.
@@
VAR 22.07 14.04 31.02

### v2

@@ 5.5.3 Types of models
As for the proposed models, 35 out of 66 belong to the pure convolutional GNN paradigm, 12 to hybrid convolutional-attention architectures, followed by 14 attentional GNNs, and 1 recurrent GNN. Two papers use both convolutional and attentional models, and other two (Refs. [art_438, art_451]) do not fall into a specific class. Regarding the definition of the graph structure, most of the papers discuss it in detail. Half of the papers published before 2024 are based on a pre-defined adjacency matrix, while the majority of papers of 2024 propose models that learn the adjacency matrix on their own, either from scratch or based on a pre-initialization (e.g., Refs. [art_108, art_444]). Here, more than in any other thematic group, there are multiple ways to define the adjacency matrix. This can be made by taking into account the structure of the road network, the actual flow of people, or a combination of different sources of information. More in detail, a first group of papers uses road connectivity (e.g., Refs. [art_104, art_120, art_123, art_135, art_431]); second group uses spatial distance (e.g., Refs. [art_93, art_96, art_99, art_110, art_128, art_132, art_434, art_454]); a third group incorporates extra information about the so-called points of interest (such as Refs. [art_94, art_96]); a fourth group considers some similarity measures between the time series (e.g., Refs. [art_97, art_117, art_122, art_123, art_435, art_445]), such as the correlation coefficient or cosine similarity; a fifth group directly uses the number of people traveling on the road at a given time step (e.g., Refs. [art_94, art_102, art_129, art_133]), or a combination of different aspects (e.g., Refs. [art_420, art_444, art_448, art_449]).
@@
The two most common loss functions used to train the models are the MAE and the MSE. Few papers consider a combination of loss functions, especially when dealing with a model that has to learn the graph structure by itself (Refs. [art_126, art_130, art_446, art_448]). Not all papers specify the language or library used for writing the code. Among those that do that, Python is the most common language, with the majority using PyTorch, whereas only a few researchers use TensorFlow. In addition, a few papers provide a link to the source code for the proposed model. The papers that provide such links are listed in Tab. LABEL:t:link_codici_traffico.
Table 21: List of source codes of the “Mobility” models in the review.
@@
Adaptive multigraph convolutional networks (AMGCN) [art_444] https://github.com/hfimmortal/AMGCN
Automated dilated spatio-temporal synchronous graph network (Auto-DSTSGN) [art_108] https://github.com/jinguangyin/Auto-DSTSGN
BigST [art_447] https://github.com/usail-hkust/BigST
@@
Progressive Graph Convolutional Network (PGCN) [art_441] https://github.com/yuyolshin/PGCN
Scale-aware neural architecture search framework for MTS forecasting (SNAS4MTF) [art_408] https://github.com/shangzongjiang/SNAS4MTF
@@
Space-Time Separable Graph Convolutional Network (STSGCN) [art_443] https://github.com/FraLuca/STSGCN
Two-level Resolution Neural Network (TwoResNet) [art_430] https://github.com/semink/TwoResNet
Wavelet dynamic spatiotemporal aware Graph Neural Network (W-DSTAGNN) [art_419] https://github.com/yash-jakhmola/w-dstagnn
@@ 5.5.4 Benchmark models
Tab. LABEL:t:benchmark_models_traffico lists the benchmark models used within the papers in the “Mobility” thematic group. These benchmarks are categorized into mathematical and statistical methods, non-GNN machine learning models, and GNN models, and only those appearing in more than two papers are included in the table.
Table 22: List of benchmark models in the “Mobility” group divided per category.
@@
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_103], [art_106], [art_109], [art_116], [art_123], [art_126], [art_129], [art_132], [art_135], [art_406], [art_408], [art_419], [art_421], [art_424], [art_430], [art_433], [art_438], [art_444], [art_446], [art_448], [art_449], [art_451], [art_452], [art_453]
Historical average (HA) [art_93], [art_94], [art_96], [art_98], [art_100], [art_102], [art_103], [art_109], [art_110], [art_120], [art_123], [art_125], [art_131], [art_132], [art_135], [art_406], [art_408], [art_420], [art_424], [art_433], [art_435], [art_438], [art_441], [art_443], [art_446], [art_447], [art_449], [art_451], [art_452], [art_453]
Vector autoregression (VAR) [art_99], [art_100], [art_126], [art_131], [art_132], [art_406], [art_419], [art_424], [art_435], [art_443], [art_445], [art_447], [art_453]
Non-GNN machine learning methods Convolutional long-short term memory network (ConvLSTM) [ConvLSTM] [art_98], [art_446]
Deep multi-view spatial-temporal network (DMVST-Net) [DMVST-Net] [art_96], [art_451]
Feed-forward neural network (FNN) [art_93], [art_97], [art_100], [art_105], [art_114], [art_120], [art_121], [art_124], [art_126], [art_408], [art_420], [art_431], [art_438]
Gated recurrent unit (GRU) [art_98], [art_99], [art_106], [art_121], [art_123], [art_124], [3], [art_131], [art_132], [art_135], [art_425], [art_452]
Long-short term memory network (LSTM) [art_93], [art_96], [art_98], [art_99], [art_100], [art_106], [art_108], [art_109], [art_113], [art_115], [art_116], [art_120], [art_121], [art_122], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_424], [art_425], [art_430], [art_433], [art_435], [art_438], [art_441], [art_442], [art_443], [art_444], [art_445], [art_449], [art_453]
@@
Seq2Seq [Seq2seq] [art_105], [art_124], [art_451]
Spatial and temporal normalization for multivariate time series forecasting (ST-Norm) [ST-Norm] [art_116], [art_117], [art_406], [art_422]
Spatial-temporal identity (STID) [STID] [art_116], [art_422], [art_433], [art_434]
Support vector regression (SVR) [art_99], [art_100], [art_103], [art_109], [art_119], [art_123], [art_129], [art_130], [art_135], [art_408], [art_419], [art_421], [art_435], [art_444], [art_445], [art_453]
Temporal convolutional network (TCN) [TCN] [art_446], [art_453]
Temporal 2D-variation modeling for general time series analysis (TimesNet) [TimesNet] [art_434], [art_435]
Transformer [attention] [art_424], [art_454]
@@
GNN methods Adaptive graph convolutional recurrent network (AGCRN) [AGCRN] [art_98], [art_100], [art_110], [art_117], [art_124], [art_125], [art_130], [art_131], [art_406], [art_419], [art_435], [art_441], [art_447], [art_448], [art_452], [art_453], [art_454]
Attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN] [art_93], [art_98], [art_99], [art_100], [art_108], [art_110], [art_113], [art_115], [art_117], [art_121], [art_125], [art_126], [3], [art_130], [art_131], [art_132], [art_406], [art_421], [art_424], [art_427], [art_432], [art_433], [art_435], [art_442], [art_445], [art_446], [art_447], [art_448], [art_449], [art_453], [art_454]
Attention based Spatial-Temporal Graph Neural Network (ASTGNN) [ASTGNN] [art_422], [art_435], [art_446]
Automated spatio-temporal graph (AutoSTG) [AutoSTG] [art_108], [art_408]
@@
Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_93], [art_94], [art_96], [art_100], [art_103], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_122], [art_124], [art_125], [art_126], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_427], [art_430], [art_432], [art_433], [art_435], [art_438], [art_441], [art_442], [art_444], [art_445], [art_447], [art_448], [art_451], [art_452], [art_453]
Dynamic graph convolutional recurrent network (DGCRN) [art_100] [art_110], [art_406], [art_444]
Dynamic and Multi-faceted SpatioTemporal Graph Convolution Network (DMSTGCN) [DMSTGCN] [art_441], [art_444]
Dynamic spatial-temporal aware graph neural network (DSTAGNN) [DSTAGNN] [art_125], [art_419], [art_424], [art_447], [art_448]
Fully connected gated graph architecture (FC-GAGA) [FC-GAGA] [art_116], [art_408], [art_427]
@@
Graph convolutional network (GCN) [art_104], [art_135], [art_431]
Graph multi-attention network (GMAN) [GMAN] [art_100], [art_109], [art_110], [art_116], [art_122], [art_128], [art_408], [art_419], [art_422], [art_427], [art_430], [art_444], [art_451]
Graph WaveNet (GWN) [GWN] [art_93], [art_100], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_419], [art_420], [art_422], [art_424], [art_430], [art_432], [art_433], [art_434], [art_435], [art_441], [art_444], [art_447], [art_448], [art_449], [art_452], [art_453], [art_454]
Discrete graph structure learning for time series (GTS) [GTS] [art_420], [art_433]
Long short-term graph convolutional networks (LSGCN) [LSGCN] [art_99], [art_432], [art_446], [art_453]
Multi-range attentive bicomponent graph convolutional network (MRA-BGCN) [MRA-BGCN] [art_109], [art_116], [art_408]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_100], [art_109], [art_110], [art_113], [art_116], [art_117], [art_122], [art_124], [art_128], [art_406], [art_408], [art_420], [art_422], [art_444], [art_445]
@@
Spectral temporal graph neural network for multivariate time-series forecasting (StemGNN) [StemGNN] [art_406], [art_422]
Spatial-temporal fusion graph neural network (STFGNN) [STFGNN] [art_98], [art_108], [art_115], [art_117], [art_125], [art_130], [art_406], [art_435], [art_445], [art_452], [art_453]
Spatial-temporal graph to sequence model (STG2Seq) [STG2Seq] [art_99], [art_424]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_93], [art_94], [art_96], [art_99], [art_100], [art_106], [art_108], [art_109], [art_110], [art_113], [art_115], [art_116], [art_117], [art_120], [art_121], [art_122], [art_124], [art_126], [3], [art_128], [art_129], [art_130], [art_131], [art_132], [art_406], [art_408], [art_419], [art_420], [art_421], [art_422], [art_424], [art_425], [art_433], [art_435], [art_438], [art_442], [art_443], [art_444], [art_445], [art_448], [art_449], [art_451], [art_452], [art_453], [art_454]
@@
Spatial-temporal graph ODE networks (STGODE) [STGODE] [art_99], [art_108], [art_117], [art_125], [art_406], [art_424], [art_447], [art_453]
Spatio-temporal graph attention network (ST-GRAT) [ST-GRAT] [art_116], [art_444]
Deep-meta-learning based model (ST-MetaNet) [ST-MetaNet] [art_96], [art_100], [art_109], [art_116], [art_408], [art_451]
Spatio-temporal synchronous graph convolutional network (STSGCN) [STSGCN] [art_98], [art_99], [art_100], [art_108], [art_115], [art_117], [art_125], [art_130], [art_131], [art_406], [art_419], [art_424], [art_435], [art_441], [art_442], [art_443], [art_453], [art_454]
Temporal graph convolutional network (T-GCN) [T-GCN] [art_103], [art_113], [art_121], [art_123], [art_421], [art_433]
Time zigzags at graph convolutional networks for time series forecasting (Z-GCNETs) [Z-GCNETs] [art_125], [art_424], [art_441]
As for the basic mathematical and statistical methods, two widely used benchmarks are the ARIMA model and the historical average model (HA), which uses the average of past observations as the forecast for future values. Despite their apparent simplicity, these benchmarks are sometimes shown to be very powerful, with an accuracy comparable to that of more complex models. As for the non-GNN machine learning models, LSTM networks are widely recognized as a type of recurrent neural network capable of handling long-term dependencies in traffic data. As for GNN-based models instead, the three most used models are the spatio-temporal graph convolutional network (STGCN) [STGCN], the diffusion convolutional recurrent neural network (DCRNN) [DCRNN], and the attention-based spatial-temporal graph convolutional network (ASTGCN) [ASTGCN], all developed specifically for the traffic forecasting problem.
The STGCN model proposed by Yu et al. in 2018 consists of two spatio-temporal convolutional blocks (with two temporal gated convolutional layers, one spatial graph convolutional layer in the middle, and residual connections) and a fully-connected output layer in the end. In its original implementation, it processes graphs a with weighted adjacency matrix based on spatial distance between sensors in the road network. The source code for the STGCN model is available at https://github.com/VeritasYin/STGCN_IJCAI-18. The DCRNN model proposed by Li et al. in 2017 aims to capture the spatial dependency by using bidirectional random walks on the graph (whose adjacency matrix is based on node proximity), and to capture the temporal dependencies by using an encoder-decoder architecture with scheduled sampling. The code can be accessed at https://github.com/liyaguang/DCRNN. The ASTGCN model proposed by Guo et al. in 2019 models at once recent, daily-periodic and weekly-periodic dependencies as three independent components, with a spatio-temporal attention mechanism for capturing the dynamic spatio-temporal correlations and a spatio-temporal convolution which simultaneously employs graph convolutions to capture the spatial patterns. The source code for the ASTGCN model is available at https://github.com/Davidham3/ASTGCN-2019-mxnet. For a more comprehensive list of source codes for the most dated benchmark GNN models in the traffic domain see Ref. [review_traffico_jiang].
@@ 5.5.5 Results
In all papers mentioned, the authors claim that the proposed GNN-based models are able to outperform the benchmarks. However, there are a limited but interesting number of situations in which some simple statistical and mathematical models demonstrate comparable accuracy, as it can be observed in Tabs. LABEL:t:risultati_metr-la and LABEL:t:risultati_pems-bay [art_452].
For the comparison and evaluation of these models, the most commonly used error metrics are MAE, RMSE, and MAPE. A few papers in addition use further metrics such as the coefficient of determination $R^{2}$. However, despite the use of the same metrics and the availability of widely used datasets, comparison among these models is not trivial, because different papers often focus on different time windows of a dataset.
@@
Table 23: Comparison of the average accuracy of the different models on the METR-LA dataset for 15, 30 and 60 minute time horizons, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the “Generic” group. The smallest errors are underlined. Numbers from the original papers.
@@
Z-GCNETs 3.23 7.87 7.48 3.93 9.75 9.40 4.83 12.04 11.57
Clearly, the results on the METR-LA dataset are not uniform, and the best performing models vary depending on the metric and the time horizon. However, the multichannel fusion graph neural network (MCFGNN) [art_86] from the “Generic” group consistently exhibits low errors across different time horizons. The MCFGNN model, proposed by Chen et al. in 2022, is composed of three main components: a graph construction layer, a graph convolution layer, and a temporal convolution layer. The input data are first processed by the graph construction layer, which incorporates three channels to capture meta-features, time-varying patterns, and globally stabilized graphs. Spatial and temporal convolutions are then applied in pairs to capture spatio-temporal dependencies in the data. To mitigate the issue of vanishing gradients, residual connections are added to each spatio-temporal block.
@@
Table 24: Comparison of the average accuracy of the different models on the PEMS-BAY dataset for 15, 30 and 60 minute time horizons, expressed in terms of MAE, MAPE and RMSE. The smallest errors are underlined. Numbers from the original papers.
@@
SLCNN 1.44 3.00 2.90 1.72 3.90 3.81 2.03 4.80 4.53
SNAS4MTF [art_408] 1.30 2.72 2.77 1.61 3.63 3.68 1.89 4.44 4.36
STAEFormer 1.31 2.76 2.78 1.62 3.62 3.68 1.88 4.41 4.34
StemGNN 1.44 3.08 3.12 1.93 4.54 4.38 2.57 6.55 5.88
STID 1.30 2.73 2.81 1.62 3.68 3.72 1.89 4.47 4.40
STFGNN 1.42 2.92 2.97 1.67 3.92 3.94 1.98 4.83 4.74
@@
Z-GCNETs 1.36 2.88 2.86 1.68 3.79 3.78 1.98 4.60 4.53
The results on the the PEMS-BAY dataset vary depending on the metrics and the time horizon. The most accurate models are the adaptive multigraph convolutional networks (AMGCN) [art_444], the multi view spatial-temporal graph neural network (MVST-GNN) [art_109], and the temporal decomposition based dynamic multi granularity graph neural network (TD2MG2NN) [art_422]. The AMGCN model was proposed by Li et al. in 2024, and includes a graph fusion block, an adaptive graph generator block, a stack of spatio-temporal convolution blocks, and an output module. The graph fusion block integrates information from five pre-defined graph structures, based on spatial distance, reachability, correlation, and distributions. Then, an independent adaptive graph generator learns an evolving graph structure to complement the pre-defined graph. Spatio-temporal features are finally extracted through a stack of graph convolutional modules with residual connections, and a time-mixing strategy. The MVST-GNN model, proposed by Li et al. in 2022, includes a multi-view temporal transformer module and a multi-view graph convolution module. The multi-view temporal transformer module is designed to capture dynamic temporal features from recent, daily, and weekly patterns using self-attention and a fusion module. Then, the graph convolution module operates on a self-learned adjacency matrix and passes its output through two fully connected linear layers to generate the final forecast. The TD2MG2NN model, instead, was introduced by Zang et al. in 2024 for evaluating the resilience of road networks in real time during heavy rain events. The process starts by decomposing the time series into seasonal and trend components. These components are processed individually through a season-temporal encoder layer with temporal convolutions, and a trend-temporal encoder layer with regression, respectively. The model then learns two feature matrices, one for daily and one for weekly patterns, which are fed into a dynamic multi-granularity graph convolution module. This module processes the feature matrices in parallel, and then passes them through a spatial decoder layer to produce the final output.
@@
Table 25: Comparison of the average accuracy of the different models on the PeMS03 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the “Generic” group. The smallest errors are underlined. Numbers from the original papers.
@@
Table 26: Comparison of the average accuracy of the different models on the PeMS04 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the “Generic” group. The smallest errors are underlined. Numbers from the original papers.
@@
Table 27: Comparison of the average accuracy of the different models on the PeMS07 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The smallest errors are underlined. Numbers from the original papers.
@@
DCRNN 25.22 11.82 38.61
DGCRN 19.91 8.32 31.47
DMSTG 19.98 8.56 33.08
DMSTGCN 20.77 8.94 33.67
DSTAGNN 21.42 9.01 34.51
DSTCGCN 20.37 8.64 33.47
DyHSL 18.84 8.11 31.65
GMAN 21.56 9.51 34.97
@@
Table 28: Comparison of the average accuracy of the different models on the PeMS08 dataset for a 60-minute time horizon, expressed in terms of MAE, MAPE and RMSE. The results of the models marked with are taken from papers in the “Generic” group. The smallest errors are underlined. Numbers from the original papers.
@@
TimesNet 19.01 11.83 30.65
VAR 19.19 13.10 29.81
As stated above, the results consistently indicate that the periodicity aware spatial-temporal adaptive hypergraph neural network (PAHNN) [art_435] proposed by Zhao et al. in 2024 is the most accurate model. The PAHNN model includes a temporal multi-periodic to time dependencies and periodic features of traffic time series, and a spatial adaptive hypergraph neural network to capture higher-order correlations among nodes.